This Question: 1 pt 6 of 6 (0 complete) The overhead reach distances of adult females are normally distributed with a mean of 197.5 cm and a standard deviation of 7.8 a. Find the probability that an individual distance is greater than 206.80 cm. b. Find the probability that the mean for 25 randomly selected distances is greater than 196.20 cm. c. Why can the normal distribution be used in part (b), even though the sample size does not exceed 30? a. The probability is (Round to four decimal places as needed.) b. The probability is (Round to four decimal places as needed.) Click to select your answer(s) * Previous

1

The overhead reach distances of adult females are normally distributed with a mean of 197.5 cm and a standard deviation of 7.8 cm. a. Find the probability that an individual distance is greater than 206.80 cm b. Find the probability that the mean for 25 randomly selected distances is greater than 196.20 cm. c. Why can the normal distribution be used in part (b), even though the sample size does not exceed 30? O A. The normal distribution can be used because the finite population correction factor is small. O B. The normal distribution can be used because the mean is large O C. The normal distribution can be used because the original population has a normal distribution O D. The normal distribution can be used because the probability is less than 0.5 Click to select your answer(s) 4 Previous QI

2:

media%2F325%2F325d1094-c151-4c1b-95dd-8b

Submit Quiz This Quiz: 6 pts possible Question Help Next

3:

media%2Fa41%2Fa41e609f-f74b-4d09-99aa-fd

4.

Find the area of the shaded region. The graph to the right depicts IQ scores of adults, and those scores are normally distributed with a mean of 100 and a standard deviation of 15. Click to view page 1 of the table. Click to view page 2 of the table. The area of the shaded region is (Round to four decimal places as needed.) Enter your answer in the answer box. Previous Hi e

This Quiz: 6 pts possibl Question Help : 80 115 2 Next . 7 PM
Answer #1

Solution:- Given mean = 197.5, standard deviation = 7.8

a. P(X > 206.80) = P((X-μ)/σ > (206.80-197.5)/7.8)
= P(Z > 1.1923)
= 0.1170

b. For n = 25

P(X > 196.20) = P((X-μ)/(σ/sqrt(n)) > (196.20-197.5)/(7.8/sqrt(25)))
= P(Z > -0.8333)
= 0.7967

c. option C. The normal distribution can be used because the orginal population has a normal distribution.
----------------

2. Z = 0.2533

x = μ + Z*σ = 100 + (0.2533*15) = 103.7995 = 103.8 (rounded)

----------------------

3. P(X < 129) = P(Z < (129-105)/20)
= P(Z < 1.2)
= 0.8849

The probability that a randomly selected adult has an IQ less than 129 is 0.8849


-----------------------

4.

=> P(80 < X < 115) = P((80-100)/15 < Z < (115-100)/15))
= P(-1.3333 < Z < 1)
= 0.7495

The area of the shaded region is 0.7495